# Parameter estimation of range-migrating targets using OTFS signals from LEO satellites

### v2

@@ IV-A Vectorized model
where $\bm{\Psi}$ is an $(NM\times NM)$-dimensional matrix with $(\bm{\Psi})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Psi_{k^{\prime},l^{\prime}}\big[k-k^{\prime},l-l^{\prime}\big]$. Interestingly, the models in (31) and (47) present the same structure; however, the involved matrices $\bm{\Phi}$ and $\bm{\Psi}$ are different. For future reference, we also define the normalized target echo as
$\bm{e}(d,v,B)=\bm{\Psi}(d,v,B)\bm{x},$ (48)
where the dependence upon the initial range $d$ of the target, the range-rate $v$ of the target, and $B$ has been made explicit.
IV-B Approximated model
We derive here an approximated model: the goal is to obtain an input-output relationship with a structure similar to the one in (32), which will turn out useful in Sec. V-A.
@@ IV-B Approximated model
whose magnitude tends to zero for large $N/B$. Upon neglecting the above term in (40) and upon ignoring the fractional part of the target range, then (41) can be approximated as
@@
$\bm{y}\approx\alpha\big(\bm{X}\odot\bm{\Theta}\odot\bm{\Lambda}(v)\big)\bm{\phi}(d,v,B),$ (52)
where $\bm{X}$ and $\bm{\phi}(d,v,B)$ are the same quantities in (32), while $\bm{\Theta}$ and $\bm{\Lambda}(v)$ are $NM\times NM$ matrices with $(\bm{\Theta})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Theta_{k^{\prime},l^{\prime}}[k,l]$ and $\big(\bm{\Lambda}(v)\big)_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{NM}l}=\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$.
To proceed, notice that $\{\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}\}_{l=-\infty}^{\infty}$ is a complex exponential whose numerical frequency depends upon the target range-rate. Since $\frac{v_{\max}T}{\lambda}<1$ and the samples $\{\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}\}_{l=0}^{M-1}$ present an overall phase variation less than $2\pi$, we propose to approximate $\bm{\Lambda}(v)$ by $\bm{\Lambda}(v_{\max}/2)$. Hence, we finally obtain
$\bm{y}\approx\alpha\bm{\Xi}\bm{\phi}(d,v,B),$ (53)
where $\bm{\Xi}=\bm{X}\odot\bm{\Theta}\odot\bm{\Lambda}(v_{\max}/2)$. By comparing (32) with (53), it is seen that the symbol matrix $\bm{X}$ in (32) is now replaced in (52) by the modified symbol matrix $\bm{\Xi}$ to account for the presence of rectangular filters. To compare (47) and (53), we compute the correspoding normalized root mean square error (NRMSE), namely, $(\mathrm{E}\left[\|\bm{\Xi}\bm{\phi}(d,v,B)-\bm{\Psi}(d,v,B)\bm{x}\|^{2}\right]/\mathrm{E}\left[\|\bm{\Psi}(d,v,B)\bm{x}\|^{2}\right])^{1/2}$, where the statistical expectation is with respect to $\bm{x}$, $d$, and $v$. For the system setup considered in Sec. VI, we found an NRMSE of about $0.07$ for any $B$, thus confirming the goodness of the approximated model in (53).
V Estimation of the target parameters
Upon accounting for the presence of the additive noise, after OTFS demodulation the received signal can be written as
@@ V Estimation of the target parameters
where $\bm{y}$ is the target echo and $\bm{\omega}$ is a complex circularly-symmetric Gaussian vector with covariance matrix $\sigma_{\omega}^{2}\bm{I}_{NM}$. Based on $\bm{z}$ and on the echo model $\bm{y}=\alpha\bm{e}(d,v,B)$ provided in Secs. III-B and IV-A, the radar receiver is faced with the problem of estimating the unknown target parameters, namely, $d$, $v$, and $\alpha$. We resort here to an ML estimation procedure; for a given $B$, the corresponding estimates are [45]
$\big[\hat{d}(B);\,\hat{v}(B)\big]$ $=\operatorname*{\arg\max}_{\begin{subarray}{c}[d;\,v]\in\mathcal{S}\end{subarray}}\frac{\big|\bm{e}(d,v,B)^{\mathsf{H}}\bm{z}\big|^{2}}{\|\bm{e}(d,v,B)\|^{2}},$ (55a)
$\hat{\alpha}(B)$ $=\frac{\bm{e}\big(\hat{d}(B),\hat{v}(B),B\big)^{\mathsf{H}}\bm{z}}{\big\|\bm{e}\big(\hat{d}(B),\hat{v}(B),B\big)\big\|^{2}},$ (55b)
where $\mathcal{S}=[0,r_{\max}]\times[0,v_{\max}]$. In practice, the continuous set $\mathcal{S}$ is replaced by a uniform rectangular grid $\mathcal{G}$ containing $G_{\rm r}=\lfloor r_{\max}/\Sigma_{\rm r}\rfloor$ and $G_{\rm rr}=\lfloor v_{\max}/\Sigma_{\rm rr}\rfloor$ points along the range and range-rate dimensions, respectively, where $\Sigma_{\rm r}\leq R_{\rm r}$ and $\Sigma_{\rm rr}\leq R_{\rm rr}$ are the corresponding point spacing. Since the cardinality of $\mathcal{G}$ and the size of $\bm{z}$ scale with $MN$, the computation of (55a) may become demanding for large $M$ and $N$, which is often the case in OTFS-based systems. To overcome this problem, we propose a two-step estimation procedure which operates as follows.
  1. 1.
    As shown in Sec. V-A, we first exploit the signal sparsity to compute a coarse estimate of the target response and, as a by-product, coarse estimates of the initial range and range-rate, denoted as $\tilde{d}(B)$ and $\tilde{v}(B)$, respectively.
  2. 2.
    Then, we restrict the search set in (55a) to a neighborhood of the above coarse estimates; in particular, we consider a uniform rectangular grid $\tilde{\mathcal{G}}$ that is centered at $\big[\tilde{d}(B);\,\tilde{v}(B)\big]$ and has $\tilde{G}_{\rm r}$ and $\tilde{G}_{\rm rr}$ points along the range and the range-rate dimension spaced of $\Sigma_{\rm r}$ and $\Sigma_{\rm rr}$, respectively, with $\tilde{G}_{\rm r}\ll G_{\rm r}$ and $\tilde{G}_{\rm rr}\ll G_{\rm rr}$.
V-A Coarse estimation of the initial range and range-rate
Consider first ideal shaping filters. Then, from (32) the received signal can be expended as $\bm{z}=\bm{X}\bm{f}+\bm{\omega}$, where $\bm{f}=\alpha\bm{\phi}$ is the target response vector. [^5] For brevity, the dependence upon $d$, $v$, and $B$ is kept concealed here. To proceed, notice that $\bm{f}$ can be partitioned into $M$ sub-vectors (blocks) of length $N$, namely, $\bm{f}=(\bm{f}_{0};\cdots;\bm{f}_{M-1})$, where $\bm{f}_{m}=\big(\alpha\Phi[0,m];\ldots;\alpha\Phi[N-1,m]\big)$; according to Remark 2, $\bm{f}$ is a block sparse vector for large $N/B$ and $M$. Notice next that the size of $\bm{z}$ can be reduced by employing a $NK\times NM$ selection matrix $\bm{S}$ whose rows are formed by any $NK$ distinct rows of $\bm{I}_{NM}$; the parameter $K\in\{1,\ldots,M\}$ rules the implementation complexity of the coarse estimator, and in practice we can select $K\ll M$. The compressed measurement $\bm{z}_{s}=\bm{S}\bm{z}\in\mathbb{C}^{NK}$ can be expanded as
@@ V-A Coarse estimation of the initial range and range-rate
$\big(\tilde{\bm{f}}_{i^{(1)}}\cdots\tilde{\bm{f}}_{i^{(\ell)}}\big)=\operatorname*{\arg\max}_{(\bm{f}_{i^{(1)}}\cdots\bm{f}_{i^{(\ell)}})\in\mathbb{C}^{N\times\ell}}\Big\|\bm{z}_{s}-\sum_{q=1}^{\ell}\bm{X}_{s,i^{(q)}}\bm{f}_{i^{(q)}}\Big\|^{2};$ (58)
finally, the residual vector and the search set are updated as $\bm{z}^{(\ell+1)}_{s}=\bm{z}^{(\ell)}_{s}-\sum_{q=1}^{\ell}\bm{X}_{s,i^{(q)}}\tilde{\bm{f}}_{i^{(q)}}$ and $\mathcal{I}^{(\ell+1)}=\mathcal{I}^{(\ell)}\setminus\{i^{(\ell)}\}$, respectively, and the procedure is stopped when $\ell=K$ or $\|\bm{z}^{(\ell+1)}_{s}\|^{2}<\epsilon NK\sigma_{\omega}^{2}$, with $\epsilon$ being a positive threshold. If the procedure ends in $\tilde{\ell}$ iterations, an estimate $\tilde{\bm{f}}$ of the target response vector is obtained from the indexes $i^{(1)},\ldots,i^{(\tilde{\ell})}$ and the sub-vectors $\tilde{\bm{f}}_{i^{(1)}},\ldots,\tilde{\bm{f}}_{i^{(\tilde{\ell})}}$. At this point, the entries of $\tilde{\bm{f}}$ are arranged into the target response sequence $\tilde{F}[k,l]=(\tilde{\bm{f}})_{lN+k+1}$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, which provides an estimate of $\alpha\Phi[k,l]$; finally, coarse estimates of the initial range and range-rate of target are obtained from the location of the maximum element of $|\tilde{F}[k,l]|$, in keeping with the discussion provided in Remark 4.
@@
V-B Extension to multiple targets
The above estimation procedure can be combined with the CLEAN algorithm in [37, 38, 39, 40, 16] to handle multiple targets at low complexity. The CLEAN algorithm iteratively reconstructs one target echo at the time, after removing the interference caused by the previously-estimated target echoes from the data. More specifically, assume that $P>1$ targets are present in the inspected region and denote by $\hat{\bm{y}}_{p}(B)=\hat{\alpha}_{p}(B)\bm{e}\big(\hat{d}_{p}(B),\hat{v}_{p}(B),B\big)$ the reconstructed target echo in the $p$-th iteration, where $\hat{d}_{p}(B)$, $\hat{v}_{p}(B)$, and $\hat{\alpha}_{p}(B)$ are the estimated initial range, range-rate, and amplitude, respectively; then, the proposed two-step estimation procedure is iterated on the cleaned data $\bm{z}-\sum_{i=1}^{p-1}\hat{\bm{y}}_{i}(B)$, for $p=1,\ldots,P$.

### v3

@@ III-B1 Vectorized model
where $\bm{\Psi}\in\mathbb{C}^{NM\times NM}$ has entries $(\bm{\Psi})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Psi_{k^{\prime},l^{\prime}}\big[k-k^{\prime},l-l^{\prime}\big]$. Note that the models in (30) and (45) share the same structure; however, the matrices $\bm{\Phi}$ and $\bm{\Psi}$ differ. We will denote $\bm{\Psi}$ by $\bm{\Psi}(d,v)$ when we want to make explicit its dependence on the target initial-range and range-rate.
The input-output relationship in (35), or equivalently (45), captures the combined effects of range migration, ICI, and ISI. We next introduce a more tractable approximated model that preserves the dominant delay-Doppler spreading effects induced by range migration while simplifying the dependence of the target response on the transmitted-symbol location.
III-B2 Approximated model
@@
whose magnitude tends to zero as $N/B$ increases. Upon neglecting the above term in (38) and ignoring the fractional part of the target range, then (39) is approximated as
@@
$\bm{y}\approx\alpha\big(\bm{X}\odot\bm{\Theta}\odot\bm{\Lambda}(v)\big)\bm{\phi}(d,v),$ (49)
where $\bm{X}$ and $\bm{\phi}(d,v)$ are the same quantities in (31), while $\bm{\Theta}$ and $\bm{\Lambda}(v)$ are $NM\times NM$ matrices with $(\bm{\Theta})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Theta_{k^{\prime},l^{\prime}}[k,l]$ and $\big(\bm{\Lambda}(v)\big)_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{NM}l}=\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$. To proceed, notice that $\{\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}\}_{l=-\infty}^{\infty}$ is a complex exponential whose numerical frequency depends upon the target range-rate. Since $\frac{v_{\max}T}{\lambda}<1$ and the samples $\{\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda M}l}\}_{l=0}^{M-1}$ present an overall phase variation less than $2\pi$, we propose to approximate $\bm{\Lambda}(v)$ by $\bm{\Lambda}(v_{\max}/2)$. Hence, we finally obtain
$\bm{y}\approx\alpha\bm{\Xi}\bm{\phi}(d,v),$ (50)
where $\bm{\Xi}=\bm{X}\odot\bm{\Theta}\odot\bm{\Lambda}(v_{\max}/2)$. By comparing (31) with (50), it is seen that the symbol matrix $\bm{X}$ in (31) is now replaced in (49) by the modified symbol matrix $\bm{\Xi}$ to account for the presence of rectangular shaping filters.
To compare (45) and (50), we compute the corresponding normalized root mean square error (NRMSE), namely, $(\mathrm{E}\left[\|\bm{\Xi}\bm{\phi}(d,v)-\bm{\Psi}(d,v)\bm{x}\|^{2}\right]/\mathrm{E}\left[\|\bm{\Psi}(d,v)\bm{x}\|^{2}\right])^{1/2}$, where the statistical expectation is with respect to $\bm{x}$, $d$, and $v$. For the system setup considered in Sec. V, we found an NRMSE of about $0.07$ for any choice of the receiver design parameter $B$, thus confirming the goodness of the approximated model in (50).
IV Estimation of the target parameters
Upon accounting for the additive noise, after OTFS demodulation the received signal can be written as
@@ IV Estimation of the target parameters
where $\bm{\omega}$ is a circularly symmetric complex Gaussian vector with covariance matrix $\sigma_{\omega}^{2}\bm{I}_{NM}$. Based on $\bm{z}$ and the target echo model derived in Sec. III, the radar receiver is faced with the problem of estimating the unknown target parameters, namely $d$, $v$, and $\alpha$. The corresponding ML estimates are [40]
[…]
$\hat{\alpha}$ $=\frac{\bm{e}\big(\hat{d},\hat{v}\big)^{\mathsf{H}}\bm{z}}{\big\|\bm{e}\big(\hat{d},\hat{v}\big)\big\|^{2}},$ (52b)
where $\mathcal{S}=[0,r_{\max}]\times[0,v_{\max}]$, and
$\bm{e}(d,v)=\begin{cases}\bm{\Phi}(d,v)\bm{x},&\text{for ideal shaping filters},\\ \bm{\Psi}(d,v)\bm{x},&\text{for rectangular shaping filters}.\end{cases}$ (53)
In practice, the continuous set $\mathcal{S}$ is replaced by a uniform rectangular grid $\mathcal{G}$ containing $G_{\rm r}=\lfloor r_{\max}/\Sigma_{\rm r}\rfloor$ and $G_{\rm rr}=\lfloor v_{\max}/\Sigma_{\rm rr}\rfloor$ points along the initial-range and range-rate domains, respectively, where $\Sigma_{\rm r}\leq R_{\rm r}$ and $\Sigma_{\rm rr}\leq R_{\rm rr}$ are the corresponding point spacing.
A direct implementation of (52) entails a complexity of
$\mathcal{O}\big(G_{\rm r}G_{\rm rr}(MN)^{2}\big).$ (54)
In particular, the evaluation of (52) may become computationally demanding when a fine search grid is required. To alleviate this burden, we propose in Sec. IV-A a reduced-complexity two-step estimation procedure that exploits the sparse structure of the target echo. This procedure is further extended in Sec. IV-B to also handle multiple targets.
IV-A Two-step estimation
[…]
IV-A1 Coarse estimation of the initial-range and range-rate
Consider first ideal shaping filters. From (31), the received signal can be written as $\bm{z}=\bm{X}\bm{f}+\bm{\omega}$, where $\bm{f}=\alpha\bm{\phi}$ is the target response vector. For brevity, the dependence on $d$, $v$, and $B$ is omitted in this section. To proceed, $\bm{f}$ is partitioned into $M$ subvectors (blocks) of length $N$, namely, $\bm{f}=(\bm{f}_{0};\cdots;\bm{f}_{M-1})$, where $\bm{f}_{m}=\big(\alpha\Phi[0,m];\ldots;\alpha\Phi[N-1,m]\big)$. According to Remark 3, $\bm{f}$ is block-sparse for large $N/B$ and $M$. Next, the dimension of $\bm{z}$ is reduced by means of an $NK\times NM$ selection matrix $\bm{S}$ whose rows are distinct rows of $\bm{I}_{NM}$. The parameter $K\in\{1,\ldots,M\}$ controls the implementation complexity. The compressed measurement $\bm{z}_{s}=\bm{S}\bm{z}\in\mathbb{C}^{NK}$ can be expressed as
@@ IV-A1 Coarse estimation of the initial-range and range-rate
$\big(\tilde{\bm{f}}_{i^{(1)}}\cdots\tilde{\bm{f}}_{i^{(\ell)}}\big)=\operatorname*{\arg\min}_{(\bm{f}_{i^{(1)}}\cdots\bm{f}_{i^{(\ell)}})\in\mathbb{C}^{N\times\ell}}\Big\|\bm{z}_{s}-\sum_{q=1}^{\ell}\bm{X}_{s,i^{(q)}}\bm{f}_{i^{(q)}}\Big\|^{2};$ (57)
finally, the residual vector and the search set are updated as $\bm{z}^{(\ell+1)}_{s}=\bm{z}^{(\ell)}_{s}-\sum_{q=1}^{\ell}\bm{X}_{s,i^{(q)}}\tilde{\bm{f}}_{i^{(q)}}$ and $\mathcal{I}^{(\ell+1)}=\mathcal{I}^{(\ell)}\setminus\{i^{(\ell)}\}$, respectively, and the procedure is stopped when $\ell=K$ or $\|\bm{z}^{(\ell+1)}_{s}\|^{2}<\epsilon NK\sigma_{\omega}^{2}$, with $\epsilon$ being a positive threshold. If the procedure ends in $\tilde{\ell}$ iterations, an estimate $\tilde{\bm{f}}$ of the target response vector is obtained from the indexes $i^{(1)},\ldots,i^{(\tilde{\ell})}$ and the sub-vectors $\tilde{\bm{f}}_{i^{(1)}},\ldots,\tilde{\bm{f}}_{i^{(\tilde{\ell})}}$. At this point, the entries of $\tilde{\bm{f}}$ are arranged into the sequence $\tilde{F}[k,l]=(\tilde{\bm{f}})_{lN+k+1}$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, which provides an estimate of $\alpha\Phi[k,l]$; finally, coarse estimates of the target initial-range and range-rate are obtained from the location of the maximum element of $|\tilde{F}[k,l]|$.
@@
IV-A2 Implementation complexity
We now discuss the implementation complexity of the proposed two-step estimator. In the first step, the BOMP algorithm is applied to the compressed model in (55). At the $\ell$-th iteration, the block-selection requires the computation of $\bm{X}_{s,i}^{\rm H}\bm{z}_{s}^{(\ell)}$ for all remaining blocks; since $\bm{X}_{s,i}\in\mathbb{C}^{NK\times N}$, this operation has complexity $\mathcal{O}((M-\ell+1)KN^{2})$. The following least-squares update involves the matrix obtained by concatenating the $\ell$ selected blocks; a standard dense least-squares implementation therefore requires $\mathcal{O}(K\ell^{2}N^{3}+\ell^{3}N^{3})$ operations. The residual updates have lower complexity and do not affect the dominant scaling. Hence, summing over at most $K$ iterations yields a complexity $\mathcal{O}(MK^{2}N^{2}+K^{4}N^{3})$ for the first step. In the second step, the metric in (52a) is evaluated over the local grid $\tilde{\mathcal{G}}$, with complexity $\mathcal{O}\big(\tilde{G}_{\rm r}\tilde{G}_{\rm rr}(MN)^{2}\big)$. Therefore, the overall complexity of the proposed two-step estimator is
$\mathcal{O}\!\left(MK^{2}N^{2}+K^{4}N^{3}+\tilde{G}_{\rm r}\tilde{G}_{\rm rr}(MN)^{2}\right).$ (58)
Since in practice $K\ll M$ and $\tilde{G}_{\rm r}\tilde{G}_{\rm rr}\ll G_{\rm r}G_{\rm rr}$, this complexity is substantially lower than that in (54).
IV-B Extension to multiple targets
Multiple targets can be handled through the iterative CLEAN algorithm [32, 33, 34, 35], which repeatedly estimates and removes the dominant target echo from the received signal. Let $\bm{r}^{(1)}=\bm{z}$ denote the initial residual. At the $p$-th iteration, the proposed two-step estimator is applied to $\bm{r}^{(p)}$ to extract the dominant remaining target and estimate its initial-range $\hat{d}^{(p)}$, range-rate $\hat{v}^{(p)}$, and amplitude $\hat{\alpha}^{(p)}$. The corresponding echo is then reconstructed as $\hat{\bm{y}}^{(p)}=\hat{\alpha}^{(p)}\bm{e}\big(\hat{d}^{(p)},\hat{v}^{(p)}\big)$, and the residual is updated as $\bm{r}^{(p+1)}=\bm{r}^{(p)}-\hat{\bm{y}}^{(p)}$ to mitigate the interference caused by previously extracted targets in the next iteration $p+1$. If the number of targets is known, say $P$, the procedure is repeated exactly $P$ times. Otherwise, the iterations are stopped when $\|\bm{r}^{(p+1)}\|^{2}$ falls below a given threshold or when a maximum number of iterations is reached.
Since each iteration requires one execution of the proposed two-step estimator, with complexity given in (58), together with a residual update of complexity $\mathcal{O}(MN)$, the overall complexity scales linearly with the number of extracted targets. If $P$ targets are processed, the resulting complexity is
$\mathcal{O}\big(P\big(MK^{2}N^{2}+K^{4}N^{3}+\tilde{G}_{\rm r}\tilde{G}_{\rm rr}(MN)^{2}\big)\big).$ (59)
V Numerical analysis
To assess the system performance, we consider the parameters in Table I. For these values, the maximum variation of the target range during one OTFS frame is $\bar{v}_{\rm max}NT\approx 136$ m, which corresponds to about $3.5$ times the range resolution $R_{\rm r}$. The initial-range $\bar{d}$ and range-rate $\bar{v}$ are randomly generated within the intervals $\bar{r}_{\rm min},\,\bar{r}_{\rm max}]$ and $\bar{v}_{\rm min},\,\bar{v}_{\rm max}]$, respectively, while a Swerling I fluctuation model is adopted for the amplitude $\alpha$. The information symbols $\{X_{\rm DD}[k,l]\}$ are independently drawn from a unit-energy $4$-QAM constellation.
The entries of the observed vector $\bm{z}$ are generated according to (12), (15), and (19). The radar receiver, instead, is designed according to the model developed in Secs. III-A and III-B for ideal and rectangular shaping filters, respectively. Unless otherwise stated, we set $B=16$. In the implementation of the proposed two-step estimator, we use $\Sigma_{\rm r}=R_{\rm r}/100$, $\Sigma_{\rm rr}=R_{\rm rr}/100$, $\tilde{G}_{\rm r}=401$, $\tilde{G}_{\rm rr}=401$, $K=4P$, and $\epsilon=1$.
For future reference, we define the target SNR as
$\mathrm{SNR}=\mathrm{E}[|\alpha|^{2}]NM/\sigma_{\omega}^{2}.$ (60)
Also, let $d$, $v$, and $\alpha$ denote the true values of the initial-range, range-rate, and amplitude of the target, and let $\hat{d}$, $\hat{v}$, and $\hat{\alpha}$ be the corresponding estimates. We define the root mean square error (RMSE) of the initial-range and range-rate estimates, and the NRMSE of the amplitude estimate, as $\mathrm{RMSE}_{\rm r}=(\mathrm{E}[|d-\hat{d}|^{2}])^{1/2}$, $\mathrm{RMSE}_{\rm rr}=(\mathrm{E}[|v-\hat{v}|^{2}])^{1/2}$, and $\mathrm{NRMSE}_{\rm\alpha}=(\mathrm{E}[|\alpha-\hat{\alpha}|^{2}]/\mathrm{E}[|\alpha|^{2}])^{1/2}$, respectively. In the following, we first examine the single-target case and then the multi-target scenario.